OpenStudy (the_rebel_outlaw):

Can you please help me with this question? Select all of the equations that will produce exponential decay. y = (0.1)^x y = 2(1/4)^x y = (3)^x y = 1/2(5)^x y = (2)^x y = 3(0.4)^x

OpenStudy (anonymous):

what do u think it is

OpenStudy (welshfella):

for decay, y will decrease in value while x increases.

OpenStudy (the_rebel_outlaw):

I think it would be y = (0.1)^x y = 1/2(5)^x y = 3(0.4)^x

OpenStudy (anonymous):

thats wht i was thinking

OpenStudy (welshfella):

the number in the parentheses must be between 0 and 1 for it to be a decay

OpenStudy (welshfella):

so y = 1/2 (5)^x is not decay

OpenStudy (the_rebel_outlaw):

so it would be y = (0.1)^x y = 3(0.4)^x

OpenStudy (welshfella):

theres another one you missed

OpenStudy (the_rebel_outlaw):

and y=(2)^x

OpenStudy (welshfella):

no :- 2 is not between 0 and 1

OpenStudy (welshfella):

2^x gives growth 2^1 - = 2 2^2 = 4 etc

OpenStudy (welshfella):

there is one which has a fraction in the parentheses

OpenStudy (the_rebel_outlaw):

y=2(1/4)^x

OpenStudy (welshfella):

right plug some values of x in and you'll see when x gets bigger y will get smaller

OpenStudy (the_rebel_outlaw):

thank you so much!!!

OpenStudy (welshfella):

yw